OpenStudy (kkbrookly):

Determine whether each sequence converges or diverges If it converges, give the limit. 1) fn=(-3)^n 2) gn=(-1/3)^n

OpenStudy (perl):

these are both 'geometric' sequences the first sequence is (-3)^1 , (-3)^2 , (-3)^3 , ... = -3, 9, -27, 81, ... does this sequence converge?

OpenStudy (kkbrookly):

Since it continues to grow, yes? I'm not 100% sure on how to tell if a sequence converges or not.

OpenStudy (freckles):

the terms alternate they do not get closer to a number

OpenStudy (kkbrookly):

So, that means it diverges?

OpenStudy (freckles):

\[\text{ example } \\ \text{ Let } a_n=\frac{n-1}{n+1} \\ \lim_{n \rightarrow \infty} a_n=1 \\ \text{ so } a_n \text {converges to } 1 \]

OpenStudy (freckles):

a_n converges because the numbers in the sequence get closer to a number as our n gets larger

OpenStudy (kkbrookly):

Okay, so number one definitely wouldn't converge because it isn't constant.

OpenStudy (freckles):

\[-1,0,\frac{1}{3},\frac{2}{4}, \frac{3}{5},\frac{4}{6},\frac{5}{7},\frac{6}{8},\frac{7}{9},\frac{8}{10},...\] still a_n a_n converges because it gets closer to a particular number in this that number is 1 the sequence isn't a constant

OpenStudy (freckles):

or this paritcular sequence isn't a constant sequence but if we did have a sequence like \[b_n=5 \\ \text{ then we say} b_n \text{ converges to } 5 \]

OpenStudy (freckles):

do you think the second one you have converges or diverges?

OpenStudy (freckles):

now the numbers do alternate but do they alternate far from a certain number or close to a certain number

OpenStudy (kkbrookly):

It converges because it continues to increase, correct?

OpenStudy (kkbrookly):

Or since it is negative it would actually be decreasing.

OpenStudy (freckles):

that isn't what i'm saying...

ganeshie8 (ganeshie8):



OpenStudy (freckles):

that is a picture of the second one but i think we need to go back to the first because i think you still don't understand but first since @ganeshie8 was so kind to graph the second one

OpenStudy (freckles):

what do the dots look like they are close to as n gets large